# A Statistical Test for Comparing the Linkage and Admixture Model Based on Central Limit Theorems

### v1

@@ 5.2 Consistency
We prove Theorem 1, i.e. the consistency of the MLE. Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case. First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters given Assumption 3.1. This is a direct consequence of the mean value theorem applied to $r\mapsto|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Therefore, we first prove
$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0(*).$
Based on this, we almost immediately get the claim of Theorem 1.
Proving (*) is Lemma 5.11 whose proof is divided into three steps:
@@
    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$ Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we can infer (according to the majorant criterion)
@@
    Prove $\ell((X_{1},...,X_{M}),(q,r))$ convergences for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.9
@@
    Prove the claim ()(*). This is Lemma 5.10.
Lemma 5.9 is Lemma 7.8 in van Handel (2008), which easily can be adapted to the inhomogeneous case.
@@
There exist constants $\epsilon\in(0,1),\alpha$ so that
@@
Let Assumption 3.1 hold. Then, it holds
$\sup_{q,r\in\Theta}|\ell_{M}(q,r)-\ell(q,r)|\xrightarrow[]{M\to\infty}0.$
@@
    The details can be found in van Handel (2008). Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous. Then, the claim is a direct consequence of Lemma 5.10 and the compactness of $\Theta$.
@@
We first show that if the likelihood has a unique maximum in $(q^{0},r^{0})$, the MLE tends to this. Therefore, we calculate $0\leq\ell(q,r)-\ell(\hat{Q}^{M},\hat{R}^{M})\xrightarrow[]{M\to\infty}0$ according to Lemma 5.11. Additionally, we know that the MLE is unique, which leads to the claim. The details are described by van Handel (2008). ∎
@@ 5.3 Central Limit Results
To prove Theorem 2, we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed.
@@
$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}\left(q^{0},r^{0}\right)\right).$
@@
Let $H^{(q^{0},r^{0})}_{k}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$. First, we prove that
$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}H^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(H^{(q^{0},r^{0})}_{k}|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
is a martingale. We just have to prove that the sum is in $L^{1}.$ However, since every summand is in $L^{1},$ we immediately obtain the claim.
@@
$\sigma_{n}:=\sum_{k=1}^{n}\left(\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)\right)^{2}.$
Here, the notation does not represent that $\sigma_{n}$ depends on $q^{0},r^{0}.$ Finally, we infer the claim from this asymptotic distribution. All three constraints (3.18, 3.19 and 3.20), i.e.
@@
for $c\in\mathbb{N},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain. By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
Finally, we are ready to prove the central limit theorem.
@@
$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J_{\theta^{0}}.$ It holds
$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J_{\theta^{0}}|$
$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)d\xi$
@@
where we used dominated convergence. Then, with Proposition 5.12 and the assumption that the matrix $J_{\theta^{0}}$ is invertible to conclude the claim. ∎
@@ 6 Discussion
From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage. From a mathematical point of view, it is particularly interesting as its stationary distribution remains the same across all markers, even though the underlying Markov chain is inhomogeneous. This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007). In this work, we use this property to prove limit results for the MLE in the Admixture Model.
Specifically, we investigate the consistency and central limit theorems (CLTs) of the MLE in the Linkage Model. In doing so, we prove the uniqueness of the MLE in this model. This is an important result, as in the Admixture Model—even in the supervised setting—the MLE is sometimes non-unique (Pfaffelhuber and Rohde, 2022; Heinzel et al., 2025; Heinzel, 2025).
@@
Arguably, the most important application of the CLT is that they provide a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model. This test helps determine which model better fits a given dataset. To my knowledge, this is the first data-based model selection method for the Linkage Model, as recommended by Anderson and Burnham (2004).
@@
In this study, we only considered the supervised setting, i.e., the allele frequencies are assumed to be known. However, in the unsupervised setting, determining the number of ancestral populations is a major challenge in population genetics (Evanno et al., 2005; Wang, 2019; Pritchard et al., 2000; Raj et al., 2014; Verity and Nichols, 2016; Alexander and Lange, 2011). So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020). A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a Hidden Markov Model (HMM), such as those proposed in van Handel (2008); MacKAY (2002).

### v2

@@ 5.2 Consistency
We prove Theorem 1, i.e. the consistency of the MLE. Therefore, we follow the ideas by van Handel (2008), which are stated in the homogeneous case. First, we note that both the transition probabilities and the emission probabilities are Lipschitz continuous in the parameters.
Lemma 5.6.
Let $d_{m}\leq d_{up}$ for all $m\in\{1,...,M\}.$ There exists two constants $L_{e}\in\mathbb{R}$ and $L_{t}\in\mathbb{R}$ such that
$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})|$ $\leq L_{e}|(q,r)-(q^{\prime},r^{\prime})|$
$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$ $\leq L_{t}|(q,r)-(q^{\prime},r^{\prime})|.$
Proof.
It holds
$|\mathbb{P}_{q,r}(X_{m}=x|Z_{m}=z_{m})-\mathbb{P}_{q^{\prime},r^{\prime}}(X_{m}=x|Z_{m}=z_{m})|$ $=|q^{\prime}_{z_{m}}-q_{z_{m}}|p_{z_{m},m}.$
Hence, we can choose $L_{e}=1.$ It also holds
$|\mathbb{P}_{q,r}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})-\mathbb{P}_{q^{\prime},r^{\prime}}(Z_{m}=z_{m}|Z_{m-1}=z_{m-1})|$
$=|(1-e^{-rd_{m}})q_{k}-(1-e^{-r^{\prime}d_{m}})q^{\prime}_{k}|$
$=\big|(1-e^{-rd_{m}})(q_{k}-q^{\prime}_{k})+q^{\prime}_{k}(e^{-r^{\prime}d_{m}}-e^{-rd_{m}})\big|$
$\leq|1-e^{-rd_{m}}|\;|q_{k}-q^{\prime}_{k}|+|q^{\prime}_{k}|\,|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|$
$\leq|q_{k}-q^{\prime}_{k}|+|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|.$
Moreover, by the mean value theorem there exists $\xi\in[r,r^{\prime}]$ such that
$|e^{-rd_{m}}-e^{-r^{\prime}d_{m}}|=d_{m}\,|r-r^{\prime}|\,e^{-\xi d_{m}}\leq d_{m}\,|r-r^{\prime}|.$
We already know that the true value is the unique maximum point of the likelihood according to Theorem 4. Hence, we first prove
$\sup_{(q,r)\in\Theta}|\ell((X_{1},...,X_{M}),(q,r))-\ell(q,r)|\xrightarrow[]{M\to\infty}0$ (5)
under Assumption 3.1. Based on this, we almost immediately get the claim of Theorem 1.
Proving (5) is Lemma 5.9 whose proof is divided into three steps:
@@
    Prove $\lim_{M\to\infty}\frac{1}{M}\sum_{i=1}^{M}\mathbb{E}\left(D^{q^{0},r^{0}}_{i}\right)=\mathbb{E}\left(\ell\left(q^{0},r^{0}\right)\right).$
@@
    Prove $\ell((X_{1},...,X_{M}),(q,r))$ converges for $M\to\infty$ a.s. We call the limit $\ell(q,r).$ This is Lemma 5.7
@@
    Prove equation (5). This is Lemma 5.9.
Lemma 5.7 and Lemma 5.8 are Lemma 7.8 and Lemma 7.9 in van Handel (2008), respectively, and can easily be adapted to the inhomogeneous case.
@@
There exists a constant $\alpha$ so that
@@
Lemma 5.9 is stated for the homogeneous case as Proposition 7.5 in van Handel (2008).
@@
Let Assumption 3.1 hold. Then, equation (5) holds.
@@
  1. 1)
    Since both, the transition probabilities and the emission probabilities are uniformly bounded below, we infer (according to the majorant criterion)
@@
    The details can be found in van Handel (2008). Let us mention that we first prove that $\ell(q,r)$ is also Lipschitz continuous by using Lemma 5.6. Then, the claim is a direct consequence of Lemma 5.8 and the compactness of $\Theta$.
@@
We first show that if the likelihood has a unique maximum in $(q^{0},r^{0})$, the MLE tends to this. Therefore, we calculate $0\leq\ell(q,r)-\ell(\hat{Q}^{M},\hat{R}^{M})\xrightarrow[]{M\to\infty}0$ according to Lemma 5.9. Additionally, we know that the MLE is unique, which leads to the claim. The details are described by van Handel (2008), Theorem 7.6. ∎
@@ 5.3 Central Limit Results
To prove the asymptotic normality of the MLE (Theorem 2), we first prove that the first derivative of $\ell((X_{1},...,X_{M}),(q^{0},r^{0}))$ is asymptotically normally distributed. Based on this, we conclude the claim with similar techniques as in Hoadley (1971).
@@
$\lim_{M\to\infty}\frac{\nabla\ell((X_{1},...,X_{M}),(q^{0},r^{0}))}{\sqrt{M}}\xRightarrow{M\to\infty}\mathcal{N}\left(0,J^{-1}_{q^{0},r^{0}}\right).$
@@
The proof idea is the same as the one of Proposition 5 in (Douc and Matias, 2001). We first define
$N_{k}^{(q^{0},r^{0})}:=\nabla D^{(q^{0},r^{0})}_{k}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{k}\right)$.
Let $\mathcal{F}_{n}$ be the sigma field that is generated by $\nabla D^{(q^{0},r^{0})}_{k},k=1,...,n$. We notice that
$\left(M^{(q^{0},r^{0})}_{n}\right)_{n\in\mathbb{N}}:=\left(\sum_{k=1}^{n}N_{k}^{(q^{0},r^{0})}+\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}-\mathbb{E}\left(\nabla D^{(q^{0},r^{0})}_{n}\right)|\mathcal{F}_{n-1}\right)\right)_{n\in\mathbb{N}}$
is a martingale.
@@
$\sigma_{n}:=\sum_{k=1}^{n}\left(N_{k}^{(q^{0},r^{0})}\right)^{2}.$
Here, the notation omits that $\sigma_{n}$ depends on $q^{0},r^{0}.$
All three constraints (i.e. equations 3.18, 3.19 and 3.20 in Hall and Heyde (2014)), i.e.
@@
for $c\in\mathbb{R},$ follow directly by the uniformly boundedness of the transition and the emission probabilities of the hidden Markov chain. By using $\sigma^{(q^{0},r^{0})}_{n}/n$ convergences to the invertible matrix $J(q^{0},r^{0})$, we can directly infer the claim.
Finally, we are ready to prove the asymptotic normality of $\hat{Q}^{M},\hat{R}^{M}$.
@@
$\sqrt{M}(\hat{\theta}^{M}-\theta^{0})I_{M}(\theta^{0})=\frac{1}{\sqrt{M}}\sum_{k=1}^{M}\nabla D_{k}((X_{1},...,X_{M}),\theta^{0}).$
We prove $I_{M}(\theta^{0})\xrightarrow[]{M\to\infty}_{p}J\left(\theta^{0}\right).$ It holds
$\lim_{M\to\infty}|I_{M}\left(\theta^{0}\right)-J\left(\theta^{0}\right)|$
$=\lim_{M\to\infty}\int_{0}^{1}\frac{1}{M}\sum_{k=1}^{M}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0}+(\hat{\theta}^{M}-\theta^{0})\xi)-\mathbb{E}\left(\nabla^{2}D_{k}((X_{1},...,X_{M}),\theta^{0})\right)\right)d\xi$
@@
where we used dominated convergence. Then, with Proposition 5.10 and the assumption that the matrix $J\left(\theta^{0}\right)$ is invertible, we conclude the claim. ∎
@@ 6 Discussion
From a biological perspective, the Linkage Model is useful because it explains genetic data in a relatively simple framework while accounting for linkage. From a mathematical point of view, it is particularly interesting as the stationary distribution of the hidden Markov chain remains the same across all markers, even though the underlying Markov chain is inhomogeneous. This specific model has already been considered in the context of Markov chains (Saloff-Coste and Zúniga, 2007). In this work, we use this property to prove limit results for the MLE in the Linkage Model.
Specifically, we investigate the consistency and the asymptotic normality of the MLE in the Linkage Model. In doing so, we prove the uniqueness of the MLE in this model. This is an important result, as in the Admixture Model—even in the supervised setting—the MLE is sometimes non-unique (Pfaffelhuber and Rohde, 2022; Heinzel et al., 2025; Heinzel, 2025).
@@
Arguably, the most important application of the CLT is that it provides a theoretical foundation for a statistical test to compare the Linkage Model with the Admixture Model. This test helps determine which model better fits a given dataset. To my knowledge, this is the first data-based model selection method for the Linkage Model.
@@
In this study, we only considered the supervised setting, i.e., the allele frequencies are assumed to be known. However, in the unsupervised setting, determining the number of ancestral populations is a major challenge in population genetics (Evanno et al., 2005; Wang, 2019; Pritchard et al., 2000; Raj et al., 2014; Verity and Nichols, 2016; Alexander and Lange, 2011). So far, none of the existing methods perform well (Garcia-Erill and Albrechtsen, 2020). A promising approach for the Linkage Model—possibly also applicable to the Admixture Model—could be methods for choosing the order of a HMM, such as those proposed in van Handel (2008); Mackay (2002).